OpenStudy (anonymous):

MEDAL!!!!!! What is the equation of a parabola with vertex at (1, −3) and focus at (−2, −3)?

OpenStudy (anonymous):

I am aware of the formula of parabola and how to find the focus,directrix etc.

OpenStudy (anonymous):

@TuringTest

OpenStudy (anonymous):

@Data_LG2

OpenStudy (jdoe0001):

notice the vertex notice the focus how many units is the focus from the vertex?

OpenStudy (anonymous):

-4, 6

OpenStudy (jdoe0001):

?

OpenStudy (anonymous):

@jdoe0001 3 units

OpenStudy (imstuck):

\[(y+3)^{2}=-12(x-1)\]is what I get.

OpenStudy (jdoe0001):

 that means the parabola opens sideways, to the left

OpenStudy (anonymous):

agreed

OpenStudy (jdoe0001):

so is 3 units from the vertex to the focus, or "p" distance a horizonal parabola the left means, "y" is squared and "p" is negative thus \(\bf (y-{\color{blue}{ k}})^2=4{\color{brown}{ p}}(x-{\color{blue}{ h}})\qquad vertex \ (1,-3)\quad p=-3 \\ \quad \\ (y-{\color{blue}{ (-3)}})^2=4{\color{brown}{ -3}}(x-{\color{blue}{ 1}})\implies (y+{\color{blue}{ (3)}})^2=4{\color{brown}{ -3}}(x-{\color{blue}{ 1}})\)

OpenStudy (jdoe0001):

hmhm well so \(\bf (y-{\color{blue}{ k}})^2=4{\color{brown}{ p}}(x-{\color{blue}{ h}})\qquad vertex \ (1,-3)\quad p=-3 \\ \quad \\ (y-{\color{blue}{ (-3)}})^2=4{\color{brown}{ (-3)}}(x-{\color{blue}{ 1}})\implies (y+{\color{blue}{ (3)}})^2=12(x-{\color{blue}{ 1}})\)

OpenStudy (jdoe0001):

well. -12 dohh.. the parabola opens to the left ... so \(\bf (y-{\color{blue}{ k}})^2=4{\color{brown}{ p}}(x-{\color{blue}{ h}})\qquad vertex \ (1,-3)\quad p=-3 \\ \quad \\ (y-{\color{blue}{ (-3)}})^2=4{\color{brown}{ (-3)}}(x-{\color{blue}{ 1}})\implies (y+{\color{blue}{ (3)}})^2=-12(x-{\color{blue}{ 1}})\)

OpenStudy (jdoe0001):